### v2

# Cross-Lingual Interleaving for Speech Language Models

Abstract
Spoken Language Models (SLMs) aim to learn linguistic competence directly from speech using discrete units, widening access to Natural Language Processing (NLP) technologies for languages with limited written resources. However, progress has been largely English-centric due to scarce spoken evaluation benchmarks and training data, making cross-lingual learning difficult. We present a cross-lingual interleaving method that mixes speech tokens across languages without textual supervision. We also release an ENFR training dataset, TinyStories ($\sim\!42$k hours), together with ENFR spoken StoryCloze and TopicCloze benchmarks for cross-lingual semantic evaluation, both synthetically generated using GPT-4. On 360360M and 11B SLMs under matched training-token budgets, interleaving improves monolingual semantic accuracy, enables robust cross-lingual continuation, and strengthens cross-lingual hidden-state alignment. Taken together, these results indicate that cross-lingual interleaving is a simple, scalable route to building multilingual SLMs that understand and converse across languages. All resources are made open-source to support reproducibility [^1] HuggingFace Datasets Available at: Adel-Moumen/Spoken_TinyStories.
Index Terms—  Spoken language models, cross-lingual learning, interleaving, multilinguality
1 Introduction
Spoken Language Models (SLMs) aim to learn language directly from speech, without relying on textual supervision. This line of work emerged from ‘textless’ NLP [19], which combines discrete speech representations [23] and neural language modelling [3, 28] to capture both acoustic and linguistic regularities from audio alone. Within this framework, Generative Spoken Language Modelling (GSLM) formalises the objective by training an autoregressive LM on sequences of discrete speech units and reconstructing waveforms with a unit vocoder, with the motivation of broadening access to NLP in settings where written resources are scarce.
Despite rapid progress, the cross-lingual setting, which aims to train a single SLM over multiple languages, remains comparatively under-explored [25, 36, 13]. Two practical bottlenecks persist: (i) evaluations are predominantly English-centric, hindering measurement of cross-lingual semantic competence; and (ii) training methods that encourage sharing across languages often depend on text, for example via speech–text interleaving in Text–Speech LMs (TSLMs) [5, 25, 7, 35], which is not directly compatible with ‘textless’ SLMs. As a result, the original promise of ‘textless’ NLP learning from audio alone at scale and across languages has not yet been fully realised.
We address these limitations by introducing a cross-lingual interleaving scheme that mixes speech tokens from different languages within the same training sequence, using sentence-level alignments but no text tokens. Exposure to interleaved multilingual contexts encourages a shared representational subspace and promotes transfer, while remaining compatible with pure ‘textless’ pipelines. To make this practical and measurable, we curate bilingual, sentence-aligned spoken resources and new spoken semantic benchmarks. Our contributions are as follows:
  1. (i)
    We propose a cross-lingual interleaving strategy that alternates speech segments from different languages within a single sequence, promoting a shared representational subspace.
  2. (ii)
    We release a French–English, sentence-aligned spoken corpus derived from TinyStories [12] (approximately 4242k hours in total) together with bilingual spoken StoryCloze and TopicCloze benchmarks [22] for semantic evaluation synthetically generated using GPT-4 [1].
  3. (iii)
    We show on 360360M and 11B parameter SLMs that the method yields positive transfer to monolingual semantic tasks and strong cross-lingual performance, under a matched training-token budget. Additionally, representation analyses indicate stronger cross-lingual alignment in the hidden states.
Section 2 situates our approach with respect to SLMs, speech–text interleaving, and multilingual training. Section 3 formalises the modelling setup and the cross-lingual interleaving scheme; Sections 46 present datasets, experiments, and analyses.
2 Related Work
2.1 Speech Language Models
Early SLMs instantiate the ‘textless’ pipeline by modelling discrete speech units and decoding back to waveforms. GSLM [19] trains a decoder-only LM over units and synthesises speech with a unit-based vocoder; pGSLM [17] augments units with prosodic tokens to improve expressiveness; and dGSLM [24] introduces a dual-tower variant for two-channel dialogues. While these systems can produce locally fluent speech, long-form coherence is challenging. SpeechSSM [26] explores state-space models [14] for long-form generation, TWIST [15] shows that initialising from a pre-trained text Large Language Model (LLM) can boost semantics, and AudioLM [2] cascades semantic and acoustic token streams. More recently, Slamming [21] provides a competitive compute-efficient recipe, and Align-SLM [20] fine-tunes SLMs with direct preference optimisation. Our work stays within the speech-only paradigm but targets cross-lingual sharing through a new learning method.
2.2 Joint Speech–Text Interleaving
TSLMs extend SLMs with text to encourage cross-modal transfer. VoxtLM/SUTLM [5] jointly trains on ASR, TTS, and continuation, and identifies interleaving (switching between text and speech inside a sequence) as particularly effective for building a shared space. Spirit-LM [25] scales this recipe and emphasises interleaving as a key quality driver; SmolTolk [7] proposes architecture refinements for vocabulary expansion; and GLM4-Voice [36] leverages a text-to-token model [35] to generate audio tokens on-the-fly, sidestepping explicit aligners [29, 27]. However, these approaches depend on text and interleave within a single language, limiting direct applicability to textless, cross-lingual SLM training. We instead interleave speech tokens across languages without introducing text.
2.3 Cross-Lingual Training
Multilingual training for speech LMs is gaining traction but remains under-evaluated on spoken semantic understanding beyond English. Spirit-LM [25] includes multiple languages at scale but reports only English spoken evaluations; GLM4-Voice [36] trains on Chinese and English yet evaluates English semantics; and scheduled interleaving for speech-to-speech translation [13] interleaves monolingually with no cross-language interaction during continuation. In text, mixing languages can induce positive transfer [32]. Our method brings this intuition to the ‘textless’ speech regime by interleaving sentence-aligned speech sequences across languages, enabling both cross-lingual abilities and improved monolingual semantics under a matched training-token budget.
3 Spoken Language Modelling
The task of modelling spoken language from raw audio typically comprises three stages: (i) quantising a waveform $\mathbf{x}\in\mathbb{R}^{T}$ into a sequence of $L$ discrete speech units with $L\ll T$; (ii) training an SLM under a next-token objective on the quantised sequence; and (iii) synthesising a waveform from the discrete units via a neural decoder. We first describe speech tokenisation, then the SLM objective, and finally our cross-lingual interleaving scheme.
3.1 Speech Language Models
Speech tokenisers. Speech tokenisers transform raw waveforms into discrete representations. Let $\mathcal{X}$ denote the domain of audio samples and let a waveform be $\mathbf{x}=(x_{1},\dots,x_{T})\in\mathcal{X}^{T}$. An encoder $f:\mathbb{R}^{T}\!\rightarrow\!\mathbb{R}^{T^{\prime}\times d_{e}}$ produces a sequence of continuous features $f(\mathbf{x})=(\mathbf{v}_{1},\dots,\mathbf{v}_{T^{\prime}})$, sampled at a lower frame rate $T^{\prime}\ll T$. A quantiser $q$ maps these to discrete units $\mathbf{s}=(s_{1},\dots,s_{T^{\prime}})$ with $s_{i}\in\{1,\dots,K\}$, where $K$ is the codebook size. Common choices include $k$-means and residual vector quantisation (RVQ).
In this work we adopt Mimi [9], a neural audio codec [33, 8] implemented as a convolutional auto-encoder with RVQ in the latent space with semantic distillation from the WavLM representations [4]. Following prior evidence that early codebooks capture semantics, we only model the first (semantic) codebook during language modelling.
Speech language models. Given a sequence of $L$ speech tokens $\mathbf{s}=(s_{1},\dots,s_{L})$ over vocabulary $\mathcal{V}_{s}$, an SLM parameterises
$p_{\theta}(\mathbf{s})=\prod_{i=1}^{L}p_{\theta}\!\left(s_{i}\mid s_{<i}\right).$ (1)
Decoder-only transformers [31] are typically used as SLM and trained by minimising the autoregressive negative log-likelihood
$\mathcal{L}_{\mathrm{LM}}=-\sum_{i=1}^{L}\log p_{\theta}\!\left(s_{i}\mid s_{<i}\right).$ (2)
Each token $s_{i}$ is embedded via $E\in\mathbb{R}^{|\mathcal{V}_{s}|\times d}$ to $\mathbf{e}_{i}=E[s_{i}]$. A stack of $m$ causal transformer blocks maps $(\mathbf{e}_{1},\dots,\mathbf{e}_{L})$ to contextual states $(\mathbf{h}_{1},\dots,\mathbf{h}_{L})$, with $\mathbf{h}_{i}$ depending only on $s_{<i}$. A projection $U\in\mathbb{R}^{d\times|\mathcal{V}_{s}|}$ yields logits $\boldsymbol{\ell}_{i}=U^{\top}\mathbf{h}_{i}$ and the predictive distribution $p_{\theta}(s_{i}\!\mid\!s_{<i})=\mathrm{softmax}(\boldsymbol{\ell}_{i})$.
3.2 Cross-Lingual Interleaving Training
Standard SLM training predicts the next speech token within a monolingual sequence. Interleaving has been proposed in TSLMs to bridge text–speech modality gaps by switching at word boundaries [5, 25, 36, 6], but such methods rely on text tokens and are thus incompatible with textless SLMs; moreover, they do not readily support mixing multiple languages [32].
We formulate a cross-lingual interleaving scheme that operates directly on speech tokens by interleaving segments from multiple languages. Assuming sentence-level segmentation and cross-lingual sentence alignment, let $A^{(l)}_{j}$ denote the $j$th segment in language $l$, represented as a sequence of speech tokens. Given a set of languages $\mathcal{L}$, we sample a language label $l_{j}\in\mathcal{L}$ for each segment index $j$ and define the interleaved sequence as
$s=\big(A^{(l_{1})}_{1}\,\|\,A^{(l_{2})}_{2}\,\|\,A^{(l_{3})}_{3}\,\|\,\dots\big)$, (3)
where $\|$ denotes concatenation at sentence boundaries. [^2] Word-level switching is also possible when word-level alignments are available; unless stated otherwise, we use sentence-level switching. We then train with the standard next-token prediction objective on $s$, so that each prediction is conditioned on preceding segments that may come from different languages. Since Eq. 3 is language-agnostic, it can in principle be applied to any language for which speech tokens are available; this is consistent with observations that code-switching-style training can transfer beyond distant languages [32]. Having defined the mechanism, we next describe the sentence-aligned spoken corpora that make such training feasible (see Section 4).
4 Cross-Lingual Spoken Datasets
Table 1: Cross-Lingual TinyStories per-language statistics. Durations are computed over the audio stories in each language.
Split Lang Total (h) Avg (s) Min (s) Max (s) Train EN 20294.90 55.52 3.12 140.48 FR 21120.66 57.77 3.12 135.84 Validation EN 298.04 55.22 4.64 121.44 FR 311.64 57.74 4.08 120.64
Table 2: Cross-Lingual spoken StoryCloze and spoken TopicCloze per-language statistics. Durations are computed over the audio stories in each language.
Split Lang Total (h) Avg (s) Min (s) Max (s) StoryCloze EN 15.68 15.09 6.32 24.96 FR 16.75 16.12 5.12 26.16 TopicCloze EN 15.64 15.05 5.92 24.08 FR 16.76 16.13 6.48 27.20
Training SLMs is compute-intensive and data-hungry: effective systems require large speech corpora, while cross-lingual interleaving (see Section 3.2) additionally demands alignment. To make such training practical, we release a French–English, sentence-aligned spoken corpus derived from TinyStories [12], totalling approximately 4242k hours of speech (see Table 1 for statistics). In addition, we introduce bilingual spoken versions of StoryCloze and TopicCloze for evaluating cross-lingual semantic understanding (Table 2). We next summarise how these datasets are constructed.
4.1 Cross-Lingual TinyStories
TinyStories is a synthetic corpus of short narratives whose vocabulary targets 34-year-olds, generated by large language models to probe commonsense reasoning in text LMs. Its spoken variant has proved effective for SLM training, improving both modelling and generative metrics [6, 21]. Stories are self-contained, causally structured, and fit within a typical SLM context windows.
In order to preserve cross-lingual semantic coherence, we construct sentence-level alignments that preserve order and meaning. A high-quality MT system (GPT-4) [1] translates each sentence with full story context; the entire validation set and 600600k training sentences are translated into French while maintaining English alignment. We then synthesise speech with a multi-speaker TTS system [34] [^3] https://github.com/kyutai-labs/delayed-streams-modeling (a 1.61.6B model using delayed-streams modelling first introduced in Moshi [9]).
We subsample $500$ cross-lingual stories from the EN-FR sTinyStories validation set, for each candidate voice we synthesise EN–FR pairs with the same voice. Using cosine similarity between Speaker Verification (SV) [4] [^4] HuggingFace Model: microsoft/wavlm-base-sv embeddings, we retain voices scoring >0.90>0.90 on average across paired sentences, yielding $44$ voices (26 female, 18 male). We synthesise the final sets with these $44$ voices, reusing the same speaker across languages to promote cross-lingual speaker consistency. The sTinyStories corpus statistics are given in Table 1.
4.2 Cross-Lingual StoryCloze and TopicCloze
Assessing cross-lingual semantic abilities of SLMs is challenging due to the scarcity of spoken benchmarks beyond English [25, 36]. Inspired by StoryCloze [22] and its spoken instantiation [15], we introduce bilingual, sentence-aligned spoken variants—sSC and sTC—in French and English.
Each benchmark comprises 44k five-sentence stories: the model observes the first four sentences and must assign higher probability (log-likelihood) to the true ending than to an adversarial alternative. sSC uses semantically incompatible endings, probing causal/temporal commonsense; sTC draws negatives from different topics, probing topical coherence.
Corpus construction mirrors that in Section 4.1: sentences are translated with GPT-4 using full story context, then synthesised with the same TTS system to preserve timing and prosody. To minimise confusion from speaker variability, we use a single female speaker (the top SV performer on validation) across both languages. Dataset statistics are given in Table 2.
5 Experimental Setup
Having introduced the interleaving mechanism and the datasets that support it, we now describe our training, and evaluation protocol.
5.1 Models and training
We use the Mimi tokeniser [9] at $12.5$ Hz with vocabulary size K=2048K{=}2048 and RVQ with $32$ codebooks; only the first (semantic) codebook is modelled. SLMs are initialised from a 11B-parameter Llama 3.2 family checkpoint [10] and a 360360M-parameter Qwen2 family checkpoint [30], following TWIST [15], with a $2048$-token context window ($\approx 2.73$ minutes of speech). These initialisations are used for the 11B and 360360M models. The textual embedding tables are replaced with audio embeddings for the new tokens.
Training uses 44 H100100 ($80$ GiB) GPUs, per-GPU batch 153,600153{,}600 tokens (total 614,400614{,}400 tokens/step). The Adam [18] optimiser is used with settings of $(\beta_{1},\beta_{2})=(0.9,0.98)$, gradient clipping $1.0$, weight decay $0.1$, and a linear warm-up of $5\%$ to $5{\times}10^{-4}$ followed by linear decay. Inputs are packed by concatenating samples until the target length is reached.
Training proceeds in three stages to isolate the effect of interleaving: (1) EN-only pre-training for 5050k steps to simulate a high-resource SLM; (2) cross-lingual interleaving for 2020k steps with a language sampling probability of $0.5$ to encourage positive transfer from the high-resource language (EN) to a new language (FR); and (3) alternating monolingual fine-tuning in FR and EN for 1515k steps to enable monolingual generation in both languages. Unless noted, the 11B and 360360M models follow the same schedule. All comparisons are made with a matched training-token budget, meaning that each system sees the same number of speech tokens (and hence the same total speech duration at $12.5$ Hz). By matching the total training-token budget, we ensure that any gains arise from interleaving rather than from data volume or sampling differences.
5.2 Baselines
We compare to: (i) FR-only monolingual, trained solely on French; and (ii) EN+FR without interleaving, trained on mixed data but restricted to monolingual sequences. Both baselines follow the same protocol as mentioned before.
5.3 Datasets
For EN we use LibriHeavy [16] (5656k hours) plus our EN sTinyStories (see Section 4) split for a total of 7676k hours. For FR we use our FR sTinyStories ($\sim 21$k hours) split. Interleaving uses the full sTinyStories sentence-aligned EN–FR corpus described above.
5.4 Evaluation
sBLiMP [11] tests syntactic competence: for each grammatical/ungrammatical pair, the model should assign higher probability to the grammatical item.
sWUGGY [11] measures lexical discrimination by preferring a real word over a phonotactically similar nonce word.
sSC and sTC (see Section 4) evaluate semantic understanding by requiring preference for the true ending over an adversarial alternative.
[…]
6 Experimental Results
The results are organised around three questions: (i) does interleaving yield positive transfer to monolingual performance? (ii) does it induce cross-lingual abilities? and (iii) does it help build a shared structure of cross-lingual speech representations? The findings are reported for both 360360M and 11B SLMs.
Table 3: Cross-lingual results (accuracy, %). Best within each sub-block in bold, second-best underlined. Interleaving shows cross-lingual transfer in monolingual generation and cross-lingual performances while preserving semantic, syntactic, and lexical SLMs performances. The EN+FR baseline displays limited cross-lingual abilities, despite showing positive transfer.
Training setup # Parameters Token budget StoryCloze (sSC) TopicCloze (sTC) ZeroSpeech
[…]
random 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00
Baseline EN 1B 46.08 56.06 66.43 61.96 69.92
Baseline EN 360M 46.08 56.38 65.20 61.62 69.32
Baseline FR 1B 15.36 55.31 67.07
Baseline FR 360M 15.36 56.44 69.85
Baseline EN+FR 1B 61.44 15.36 55.79 57.83 52.32 50.77 66.86 71.24 57.93 58.36 62.29 62.24
Baseline EN+FR 360M 61.44 15.36 55.26 57.93 50.56 51.25 66.00 69.48 55.58 57.34 61.17 67.71
Cross-lingual Interleaving 1B 52.22 6.14 54.40 55.47 54.56 52.64 62.26 63.17 63.28 63.44 52.73 56.74
Cross-lingual Interleaving 360M 52.22 6.14 55.90 57.08 56.44 55.37 64.00 68.67 65.20 65.84 55.35 59.56
Cross-Lingual Interleaving + EN+FR Fine-Tuning 1B 61.44 15.36 55.63 58.31 55.21 55.05 67.45 70.39 62.90 63.35 61.75 69.15
Cross-Lingual Interleaving + EN+FR Fine-Tuning 360M 61.44 15.36 55.74 57.50 55.10 53.92 67.07 70.55 59.86 62.28 61.08 68.62
6.1 Cross-lingual interleaving helps monolingual capabilities
Starting from monolingual baselines in Table 3, French-only training already acquires non-trivial semantic competence with scores 56.4456.44% and 69.8569.85% on sSC and sTC for the 360360M model, respectively. The 11B model shows the same level of competence. Interestingly, in this data regime the 360360M model often matches or exceeds the 11B model on French, a pattern consistent with a capacity–data mismatch: with limited French hours, the smaller model is easier to optimise and less prone to overfitting or optimisation instabilities. Relative to English baselines, the sSC task tends to be slightly more favourable to French, while sTC is generally easier overall and, for French, yields an average increase of $2.64$ points over English.
Introducing interleaving without a final stabilisation phase produces a predictable trade-off. Semantic scores in English decrease by a small margin (e.g., for the 11B model, from 56.0656.06% to 54.4054.40%.), whereas syntactic and lexical probes (sBLiMP/sWUGGY) drop more noticeably (from 61.9661.96% and 69.9269.92% to 52.7352.73% and 56.7456.74%). An intuitive explanation is that exposure to rapidly alternating languages perturbs low-level regularities (agreement, morphology, phonotactics) more than high-level semantics. Despite this, French sSC improves modestly even though the model sees substantially fewer French tokens than the monolingual baseline; in this regime, the 360360M model reaches 57.0857.08% while the baseline with $3{\times}$ more tokens obtains 56.4456.44%. The gains therefore cannot be attributed to data volume and are better explained as transfer from English into French. TopicCloze in French may dip at this stage, suggesting that, on balance, raw interleaving trades a little monolingual polish for multilingual competence.
A brief bilingual stabilisation phase (fine-tuning on EN+FR without interleaving) restores most of the lost ground in English and consolidates the transfer into French. After stabilisation, English performance typically returns to within a hair’s breadth of the English-only baseline, particularly on lexical discrimination. Indeed, for the 11B model, sBLiMP and sWUGGY recover to 61.7561.75% and 69.1569.15%, aligning with the EN baseline. At an equivalent French token budget, the stabilisation stage surpasses its monolingual FR baseline by a few points on both sSC and sTC, achieving 58.3158.31% and 70.3970.39% versus 55.3155.31% and 67.0767.07%. This two-stage recipe consistently benefits both model sizes.
6.2 Cross-lingual abilities emerge with interleaving
[…]
Finally, we measured layer-wise cosine similarity between hidden states for 1,000 aligned EN–FR sentences and found that interleaving leads to consistently stronger cross-lingual alignment than training on mixed data without interleaving, with average cosine similarities of 0.73 (EN+FR baseline), 0.75 (interleaving), and 0.76 (interleaving+stabilisation) for the 1B models.
7 Conclusions
We introduce a new cross-lingual interleaving scheme that mixes speech tokens across sentence-aligned languages without textual supervision, and we release an ENFR training dataset, TinyStories ($\sim\!42$k), together with spoken StoryCloze and TopicCloze benchmarks for ENFR semantic evaluation. Under a matched training-token budget on 360360M11B SLMs, the method improves monolingual semantic accuracy, enables robust cross-lingual continuation, and strengthens cross-lingual hidden-state alignment. Because the method places minimal constraints on optimisation dynamics, it is well suited to scaling multilingual SLMs, thereby enabling models to understand and converse across languages.